Julio invested some money in an account that pays 9% interest and the rest in an account that pays 5%. He invested 4 times as much in the 9% account as in the 5% account. If the combined interest from both investments was $1435, how much did he invest in each?
Answer by mananth(16949)

x-------------- 5.00% per annum ------------- x
y------------- 9.00% per annum ------------y =4x
x+y =800....1
Interest----- $1435.00

5.00% x + 9.00% * 4 x = $1,435.00
0.05 x + 0.09 * 4 x = $1,435.00
0.05 x + 0.36 x = $1,435.00
0.41 x = $1,435.00
/ 0.41
x= $3,500.00 investment @ 5%
4 x = $14,000.00 investment @ 9%

CHECK
$175.00 + $1,260.00 = $1,435.00